# Application Programming Interface CTU

### v1

@@ III-C Deterministic Yaw Control
A key factor influencing performance in MPPI is the density of the simulated rollout trajectories in the space of all possible rollout trajectories. To increase this density without adding more rollouts, a dimension can be removed from the space.
This can be achieved by controlling the UAV around the yaw axis using a deterministic method. A proportional controller has been chosen for simplicity. It follows the reference heading $\bm{\mathbf{h}}^{\text{ref}}$ by computing
@@
where $k_{\text{MPPI},z}$ is the proportional feedback gain.
@@ III-D Geometric Control in Candidate MPPI Trajectories
In existing implementations, the rollout command sequences $\bm{\mathbf{u}}^{k}_{j}$ in MPPI rollouts are generated solely by adding noise $\delta{\bm{\mathbf{u}}_{j}^{k}}$ sampled from a normal distribution to the nominal command sequence $\bm{\mathbf{u}}^{\text{nom}}_{j}$. These rollouts can explore the space for a collision-free trajectory, but are not efficient at precisely following reference trajectories. Unlike rollouts generated randomly, utilizing a method controlling the UAV directly on the SE(3) group [10] allows generating rollouts that track the reference trajectory as well as a standalone SE(3) controller.
To the best of our knowledge, incorporating an SE(3) controller into the MPPI framework has not been done before. A version of the SE(3) controller [10] is incorporated and modified to output total thrust $F^{k}_{t,j}$ and desired body rates $\bm{\mathbf{\omega}}^{k}_{c,j}$ for time step $j$ of the rollout at index $k$. This effectively creates a predictive SE(3) controller and as shown in Sec. IV-B, it greatly helps with the tracking precision, especially for slower and smoother trajectories. The geometric controller is capable of following reference trajectories, but requires tuning and the ideal configuration depends on the specific UAV and trajectory to be followed. Running multiple rollouts with commands generated using geometric control enables using different parameters in each rollout. This is equivalent to testing multiple configurations of the geometric controller at once. The forward simulation of GMPPI is run in parallel on a GPU. To prevent warp divergence, the number of rollouts using the geometric controller $K_{\text{SE3}}$ must be a multiple of 32. Therefore, $K_{\text{SE3}}=32$ was chosen, with the split between random and geometric rollouts illustrated in Fig. 4.
@@ III-E Cost Function
MPPI provides a nominal trajectory as a weighted average of rollout trajectories (where higher weights $w_{k}$ are assigned to rollout trajectories with a better, i.e., lower cost behaviour (3)). Lower cost $C^{k}$ is assigned to trajectories conforming the objective of the control, while higher costs are assigned to trajectories that exhibit unwanted characteristics such as excessive maneuvers or a large tracking error. The proposed controller has three goals, which are precise trajectory tracking, smooth and stable flight, and obstacle avoidance, which must always take precedence over the remaining objectives.
@@ III-E1 Tracking Precision
To enable precise trajectory tracking, the errors in position, velocity, orientation, and angular velocity relative to their respective reference values are computed as
@@
They are multiplied by the position $c_{j}^{p}$, velocity $c_{j}^{v}$, orientation $c_{j}^{q}$ and angular velocity $c_{j}^{\omega}$ coefficients respectively. The orientation difference approximation [9] is defined as
@@
The coefficients can vary with the time index $j$ in the rollout. This allows assigning a lower cost to those rollout trajectories that deviate from the reference trajectory only in later steps of the rollout simulation, which promotes exploration and reallocates precision emphasis toward earlier timesteps of the simulated rollout trajectories. This increases trajectory tracking precision, as well as smoothness. Using the Euclidean norm for position instead of its square requires computing a square root but prevents the creation of an erratic nominal trajectory in cases where a larger deviation from the reference trajectory is necessary to avoid an obstacle.
@@ III-E2 Smoothness
To promote smooth flight, two cost components are added. First, cost is added for jerk $\norm{\bm{\mathbf{\dot{a}}}}$, which is larger than a multiple $t_{j}$ of the jerk present in the reference trajectory $\norm{\bm{\mathbf{\dot{a}}}^{\text{ref}}}$. The excess jerk quantity
$e_{j}=\text{max}\left(\norm{\bm{\mathbf{\dot{a}}}^{k}_{j}}-t_{j}\norm{\bm{\mathbf{\dot{a}}}^{\text{ref}}},0\right)$ (14)
is introduced and multiplied by a coefficient $c_{j}^{j}$. This punishes any excess control input based on how much it disturbs the smoothness of the flight, while still allowing sharp trajectories to be followed precisely. The reference and rollout jerk values are calculated using a finite-difference approximation from acceleration data. Noise amplification concerns do not apply, as both the reference and rollout trajectories are noise-free.
@@ III-E4 Resulting Cost Function
$C^{k}=\sum_{j=1}^{N}C^{k}_{j}.$ (18)
The entire GMPPI algorithm is summarized in Alg. 1. On lines 1–3, nominal and reference commands resampled using the dynamic timesteps are resampled. The algorithm then creates $K$ rollouts in parallel. On lines 4–11, $K_{\text{SE3}}$ geometric rollouts are created by selecting the SE(3) parameters to use and then running the simulation. On lines 12–16, random rollouts with proportional yaw control are calculated. All rollouts are assigned a cost on lines 17 and 18 and finally, the new nominal command is created as a weighed average of rollouts on line 20. The first nominal command is returned to be applied to the UAV.
@@ 26 $C^{k}=\sum_{j=0}^{N}C^{k}_{j}$ ;
28;
29 $\bm{\mathbf{u^{\text{nom}}}}$ = AverageWeighedCommands() // Eq. (4) ;
30
31return $\bm{\mathbf{u}}^{\text{nom}}_{0}$
@@ IV Results
All simulated experiments were conducted on a laptop with an Intel Core i7-1165G7 CPU and an NVIDIA GeForce MX450 GPU, which the proposed controller utilizes for parallel rollout simulation. Real-world experiments utilized a Jetson Orin NX high-level computing unit with an 8-core Arm Cortex-A78AE CPU and an Ampere GPU with GMPPI controller outputs fed into a PX4 low-level flight controller. All reference trajectories were generated either by an MPC-based tracker [33] or by a polynomial trajectory planner [34].
@@ IV-A UAV & Controller Parameters
The UAV mass $m$, the arm length $l$, the inertia matrix $\mathbf{J}$ and the rotor torque constant $c_{tf}$ of the simulated UAV are shown in Table I. The parameters of the UAV used for real-world experiments were similar. The GMPPI parameters along with control input limits for both the minimum and maximum thrust $F_{t,\text{min}},F_{t,\text{max}}$ as well as the maximum absolute angular rate $\omega_{\text{max}}$ are shown in Table III.
@@
For obstacle avoidance in simulation, a stereo depth camera with a range of $s=13~$\mathrm{m}$\$} is mounted to the UAV, although the rollout length is limited to $10\text{\,}\mathrm{m}$. The controller treats every visible obstacle as occupying $d_{a}=$2.0\text{\,}\mathrm{m}$\$} along the viewing ray before free space is presumed. All experiments had the camera tilted up by a fixed angle to allow full use of the sensor range. Parameters are shown in Table II.
TABLE II: Relationship of Speed and Camera Tilt
[…]
Tilt $\text{deg}]$ 8 10 16 22 22 27 27 30
TABLE III: Controller Parameters and Limits
@@ IV-B Reference Tracking
To test the ability of the GMPPI controller to follow reference trajectories, three trajectories were generated: a hover trajectory and two agile trajectories shown in Fig. 6. All trajectories specify the heading of the UAV to be equal to the projection of the velocity vectors to the horizontal plane to allow collision avoidance using front-mounted stereo camera.
@@
The results are shown in Table IV. The proposed GMPPI controller shows on average a 31% reduction in position Root Mean Square Error (RMSE) over the existing MPPI implementation [9] for agile trajectories. In the ”Hover“ trajectory, a 97% and 98% reduction of maximum velocity and acceleration respectively has been achieved. GMPPI exhibits only a 20% higher position RMSE than the standalone SE(3) controller, while being capable of obstacle avoidance. The RMSE of heading has been reduced compared to the previous MPPI implementation [9] by an average of almost 88%. These improvements are driven by features presented in Sec. III. In particular, the integration of geometric rollouts significantly improves performance across all tested trajectories, while the dynamic time step length and changing parameters improve hovering stability at the cost of slightly increased position RMSE in agile trajectories.
@@ IV-C Obstacle Avoidance
Figure 8: Graph of success rates for different controllers in the Poisson forest. Bubble results are from [11], Learning results are from [6] and Dif. physics results are from [7].
The results of the experiments are shown in Fig. 8, where the “Bubble” results are from [11], the “Learning” results are from [6], and the “Dif. physics” results are from [7]. In the given forest density, the proposed GMPPI controller achieves a 100% success rate up to a speed of 11 $1\text{\,}\mathrm{m}\text{\,}{\mathrm{s}}^{-1}$, surpassing all learning-based methods, with failures at higher speeds occurring at challenging sections of the forest with higher local densities. At $13\text{\,}\mathrm{m}\text{\,}{\mathrm{s}}^{-1}$, the proposed method has 80% less unsuccessful attempts than the Bubble planner [11]. It is worth noting that the Bubble planner uses a sensor with a range of $8\text{\,}\mathrm{m}$, while the GMPPI controller had the rollouts length limited to $10\text{\,}\mathrm{m}$.
@@ IV-D Real-world experiments
Figure 9: Real-world flight using the GMPPI controller. The dashed green line is the reference trajectory supplied by the MPC-based tracker and the filled line is the actual trajectory taken by the UAV. Obstacles are represented by black dots and the green and red dots are the start and end points respectively.
Experiments with collision-free trajectories at speeds of up to $17\text{\,}\mathrm{m}\text{\,}{\mathrm{s}}^{-1}$ and accelerations of up to $35\text{\,}\mathrm{m}\text{\,}{\mathrm{s}}^{-2}$ were conducted to validate controller performance. In a figure 8-shaped trajectory with maximum speed of $14\text{\,}\mathrm{m}\text{\,}{\mathrm{s}}^{-1}$ and acceleration of $26\text{\,}\mathrm{m}\text{\,}{\mathrm{s}}^{-2}$, GMPPI achieved a position RMSE of $0.69\text{\,}\mathrm{m}$, demonstrating no performance degradation compared to simulation. An artificially constructed forest shown in Fig. 1 was used for real-world obstacle avoidance testing to preserve localisation capability based on RTK GPS. Flight experiments were conducted at speeds ranging from $2\text{\,}\mathrm{m}\text{\,}{\mathrm{s}}^{-1}$ to $10\text{\,}\mathrm{m}\text{\,}{\mathrm{s}}^{-1}$ and the UAV successfully avoided obstacles even when following a trajectory that was planned to be colliding. An example flight path is shown in Fig. 9.
@@ V Conclusion
[…]

### v2

@@ III-C Deterministic Yaw Control
A key factor influencing performance of MPPI-based control is the density of the simulated rollout trajectories in the space of all possible rollout trajectories. To increase this density without adding more rollouts, a dimension can be removed from the space.
This can be achieved by modifying the existing sampling-based MPPI [7] rollout generation method to control the yaw axis rotation deterministically. A proportional controller has been chosen for simplicity. It follows the reference heading $\bm{\mathbf{h}}^{\text{ref}}$ by computing
@@
where $k_{\text{MPPI},z}$ is the proportional feedback gain and $\angle(\cdot,\cdot)$ denotes the angle between two vectors.
@@ III-D Geometric Control in Candidate MPPI Trajectories
In MPPI, the rollout command sequences $\bm{\mathbf{u}}^{k}_{j}$ are generated by adding random noise $\delta{\bm{\mathbf{u}}_{j}^{k}}$ to the nominal command sequence $\bm{\mathbf{u}}^{\text{nom}}_{j}$. Such rollouts explore the space for a collision-free trajectory, but are not efficient at precisely following reference trajectories. This motivates the search for rollout methods, which would allow precise trajectory tracking. One such option is a geometric SE(3) controller.
The geometric SE(3) controller [8] is a theoretically proven and practically validated method capable of following reference trajectories precisely and in a computationally efficient manner. One of the practical downsides of this method is the tuning process. It requires precise selection of gains, which often need to vary not just based on the controlled UAV, but also the specific trajectory to be followed.
MPPI and SE(3) control methods have complementary strengths and weaknesses. Using the SE(3) controller as a rollout generation method produces rollouts capable of tracking a reference trajectory while the generation remains computationally feasible. The remaining sampling-based rollouts can then focus on obstacle avoidance and exploration, as they no longer need to be used for following the reference trajectory.
Moreover, because MPPI already runs hundreds of parallel rollouts, more than a single SE(3) rollout may be added. The amount of rollouts generated using SE(3) is labelled $K_{\text{SE3}}$, with the other $K-K_{\text{SE3}}$ remaining as before. In practice, rollouts run on a GPU in warps of $32$, so $K_{\text{SE3}}=32$ is chosen, with the split between random and geometric rollouts illustrated in Fig. 4. Each of the SE(3) rollouts uses different parameters and multiple SE(3) configurations are therefore tested simultaneously, with the best configuration being selected by the proposed GMPPI cost function. Little benefit would arise from increasing or tuning the $K_{\text{SE3}}$ parameter, as the majority of the rollouts must remain sampling-based to preserve the ability to avoid obstacles. The resulting controller can precisely track reference trajectories and avoid small unmapped obstacles.
@@ III-E Cost Function
MPPI computes a nominal trajectory as a weighted average of rollout trajectories (where higher weights $w_{k}$ are assigned to rollout trajectories with a better, i.e. lower cost, behaviour (3)). Lower cost $C^{k}$ is assigned to trajectories conforming the objective of the control, while higher costs are assigned to trajectories that exhibit unwanted characteristics such as excessive maneuvers or a large tracking error. The proposed controller has three goals, which are precise trajectory tracking, smooth flight, and obstacle avoidance, which always takes precedence. In GMPPI, the cost function selects whether the geometric or sampling-based rollouts should be prioritized.
@@ III-E1 Tracking Precision
To enable precise trajectory tracking, the errors in position, velocity, orientation, and angular velocity relative to their reference values are computed as
@@
where $d_{\bm{\mathbf{q}}}(\bm{\mathbf{q}}_{1},\bm{\mathbf{q}}_{2})$ is an approximation of orientation difference adopted from [7] and defined as
@@
The errors are multiplied by the position $c_{j}^{p}$, velocity $c_{j}^{v}$, orientation $c_{j}^{q}$ and angular velocity $c_{j}^{\omega}$ coefficients respectively. The coefficients can vary with the time index $j$ in the rollout. This allows assigning a lower cost to those rollout trajectories that deviate from the reference trajectory only in later steps of the rollout simulation, which promotes exploration and reallocates precision emphasis toward earlier timesteps of the simulated rollout trajectories, allowing the controller to maximize the use of the precision of the SE(3) rollouts. This increases trajectory tracking precision, as well as smoothness. Using the Euclidean norm for position instead of its square requires computing a square root but prevents the creation of an erratic nominal trajectory in cases where a larger deviation from the reference trajectory is necessary to avoid an obstacle.
@@ III-E2 Smoothness
To promote smooth flight, two cost components are added. First, cost is added for jerk $\norm{\bm{\mathbf{\dot{a}}}}$, which is larger than a multiple $t_{j}$ of the jerk present in the reference trajectory $\bm{\mathbf{\dot{a}}}^{\text{ref}}||$. The excess jerk quantity
[…]
is introduced and multiplied by a coefficient $c_{j}^{j}$. This punishes any excess control input based on how much it disturbs the smoothness of the flight, while still allowing sharp trajectories to be followed precisely. This is a key feature allowing the controller to use the full potential of agile SE(3)-generated rollout trajectories if the reference trajectory contains sharp maneuvers. The reference and rollout jerk values are calculated using a finite-difference approximation from acceleration data. Noise amplification concerns do not apply, as both the reference and rollout trajectories are noise-free.
@@ III-E4 Resulting Cost Function
$C^{k}=\sum\nolimits_{j=1}^{N}C^{k}_{j}.$ (18)
The entire GMPPI algorithm is summarized in Alg. 1. On lines 1–3, nominal and reference commands are resampled using the dynamic timesteps. The algorithm then creates $K$ rollouts in parallel. On lines 4–11, $K_{\text{SE3}}$ geometric rollouts are created by selecting the SE(3) parameters to use and then running the simulation. On lines 12–16, random rollouts with proportional yaw control are calculated. All rollouts are assigned a cost on lines 17–18, and finally, the new nominal command is created as a weighted average of rollouts on line 19. The first nominal command is used by the UAV.
@@ 26      $C^{k}=\sum_{j=0}^{N}C^{k}_{j}$ ;
[…]
29
return $\bm{\mathbf{u}}^{\text{nom}}_{0}$
@@ IV Results
All simulated experiments were conducted on a laptop with an Intel Core i7-1165G7 CPU and an NVIDIA GeForce MX450 GPU, which the proposed controller utilizes for parallel rollout simulation. Real-world experiments utilized a Jetson Orin NX high-level computing unit with an 8-core Arm Cortex-A78AE CPU and an Ampere GPU with GMPPI controller outputs fed into a PX4 low-level flight controller. An Intel RealSense D435i depth camera set to a resolution of 640x480 px was used. All reference trajectories were generated either by an MPC-based tracker [30] or by a polynomial trajectory planner [31].
@@ IV-A UAV & Controller Parameters
The UAV mass $m$, the arm length $l$, the inertia matrix $\mathbf{J}$ and the rotor torque constant $c_{tf}$ of the simulated UAV are shown in Table I. The parameters of the UAV used for real-world experiments were similar. The GMPPI parameters along with control input limits for both the minimum and maximum thrust $F_{t,\text{min}},F_{t,\text{max}}$ as well as the maximum absolute angular rate $\omega_{\text{max}}$ are shown in Table II.
@@
For obstacle avoidance in simulation, a stereo depth camera with a range of $s=13~$\mathrm{m}$\$} is mounted to the UAV, although the rollout length is limited to $10\text{\,}\mathrm{m}$. The controller treats every visible obstacle as occupying $d_{a}=$2.0\text{\,}\mathrm{m}$\$} along the viewing ray before free space is presumed. All experiments had the camera tilted up by a fixed angle to allow full use of the sensor range. Parameters are shown in Table III.
TABLE II: Controller Parameters and Limits
@@
TABLE III: Relationship of Speed and Camera Tilt
[…]
Tilt $\text{deg}]$ 8 10 16 22 22 27 27 30
@@ IV-B Reference Tracking
To test the ability of the GMPPI controller to follow reference trajectories, three trajectories were generated: a hover trajectory and two agile trajectories shown in Fig. 6. All trajectories specify the heading of the UAV to be equal to the projection of the velocity vectors onto the horizontal plane to allow collision avoidance using a front-mounted stereo camera.
@@
The results are shown in Table IV. The proposed GMPPI controller shows on average a 31% reduction in position Root Mean Square Error (RMSE) over the existing MPPI implementation [7] for agile trajectories. In the ”Hover“ trajectory, a 97% and 98% reduction of maximum velocity and acceleration respectively has been achieved. GMPPI exhibits a 20% higher position RMSE than the standalone SE(3) controller. This is an effect introduced by the GMPPI cost function, which must consider all rollout trajectories, including the ones dedicated to exploring the space for obstacle-free areas. In return for this small decrease in precision, the ability to avoid obstacles on the controller level is gained, which is critical for avoiding the risk of controller short-cutting, especially in agile flights. The RMSE of heading has been reduced compared to the previous MPPI implementation [7] by an average of almost 88%. These improvements are driven by features presented in Sec. III. In particular, they highlight the synergy between SE(3) and MPPI control described in Sec. III-D. The combination of these methods yields an obstacle-aware controller retaining most of the precision of the standalone SE(3) controller.
@@ IV-C Obstacle Avoidance
Figure 8: Graph of success rates for different controllers in the Poisson forest in simulated experiments. Bubble results are from [9], Learning results are from [4] and Dif. physics results are from [5].
The results of the simulated experiments are shown in Fig. 8, where the “Bubble” results are from [9], the “Learning” results are from [4], and the “Dif. physics” results are from [5]. In the given forest density, the proposed GMPPI controller achieves a 100% success rate up to a speed of 11 $1\text{\,}\mathrm{m}\text{\,}{\mathrm{s}}^{-1}$, surpassing all learning-based methods, with failures at higher speeds occurring at challenging sections of the forest with higher local densities. At $13\text{\,}\mathrm{m}\text{\,}{\mathrm{s}}^{-1}$, the proposed method has 80% less unsuccessful attempts than the Bubble planner [9]. It is worth noting that the Bubble planner uses a sensor with a range of $8\text{\,}\mathrm{m}$, while the GMPPI controller had the rollouts length limited to $10\text{\,}\mathrm{m}$.
@@ IV-D Real-world experiments
Figure 9: Real-world flight using the GMPPI controller. Filled lines represent UAV trajectories; black dots denote obstacles, and green/red dots mark the start/end points. Obstacle density is $\delta=\frac{1}{12}\,\mathrm{pylons}\,\cdot\,${\mathrm{m}}^{-2}$\$ with a maximum reference velocity of $10\text{\,}\mathrm{m}\text{\,}{\mathrm{s}}^{-1}$.
TABLE V: Minimum Distance to Obstacles - Statistics
Velocity Count Mean [m] Std. Deviation [m] Worst-case [m]
$10\text{\,}\mathrm{m}\text{\,}{\mathrm{s}}^{-1}$ 15 0.38 0.14 0.18
$7\text{\,}\mathrm{m}\text{\,}{\mathrm{s}}^{-1}$ 5 0.34 0.07 0.24
$3\text{\,}\mathrm{m}\text{\,}{\mathrm{s}}^{-1}$ 6 0.56 0.09 0.44
Experiments with collision-free trajectories at speeds of up to $17\text{\,}\mathrm{m}\text{\,}{\mathrm{s}}^{-1}$ and accelerations of up to $35\text{\,}\mathrm{m}\text{\,}{\mathrm{s}}^{-2}$ were conducted to validate controller performance. In a figure 8-shaped trajectory with maximum speed of $14\text{\,}\mathrm{m}\text{\,}{\mathrm{s}}^{-1}$ and acceleration of $26\text{\,}\mathrm{m}\text{\,}{\mathrm{s}}^{-2}$, GMPPI achieved a position RMSE of $0.69\text{\,}\mathrm{m}$, showing no performance loss compared to simulation. To verify obstacle avoidance capability in the real world, a total of $42$ flights were conducted through an obstacle course composed of seven pylons, which simulated trees, but allowed preserving localisation capability based on RTK GPS. A virtual ceiling was added to prevent the UAV from flying over the obstacles. Of the $42$ conducted flights, $26$ were carried out using a straight line reference trajectory through a pylon configuration shown in Fig. 1 with a density of $\delta=\frac{1}{12}\,\mathrm{pylons}\,\cdot\,\mathrm{m}^{-2}$. These flight paths and the obstacle layout are shown in Fig. 9. The remaining experiments were carried out at speeds $3\text{\,}\mathrm{m}\text{\,}{\mathrm{s}}^{-1}$ and $7\text{\,}\mathrm{m}\text{\,}{\mathrm{s}}^{-1}$. No collision has been recorded and the statistics of minimum distances to obstacles are shown in Table V.
@@ V Conclusion
We presented GMPPI, a controller for precise agile tracking with obstacle avoidance. To our knowledge, it is the first to combine sampling-based MPPI with geometric SE(3) control. The ability to track agile trajectories was tested in simulation, with speeds up to $16~$\mathrm{m}~\mathrm{s}^{-1}$\$ and acceleration in excess of $22~$\mathrm{m}~\mathrm{s}^{-2}$\$. The results show that the controller achieves similar position RMSE to the SE(3) controller on agile trajectories. At the same time, our controller outperformed learning-based methods as well as the state-of-the-art obstacle-aware Bubble planner in obstacle avoidance capabilities. In real-world experiments, GMPPI showed no decrease in performance in following agile trajectories, and it was able to avoid obstacles at speeds of up to $10\text{\,}\mathrm{m}\text{\,}{\mathrm{s}}^{-1}$.